If the altitude of an equilateral triangle is 15, find the length of the sides.
Answer by unlockmath(1688)

Hello,
Let's have x represent the length of each side of the triangle. Now we can
write an equation as using Pythagorean formula:
x^2+1/2x^2=15^2
This can be written out as:
1 1/2 x^2=225
Divide each side by 1.5 to get:
x^2=150
Square root both sides:
x=sq rt 150 or rewritten as 5 sq rt 6
Make sense?
RJ
www.math-unlock.com